Evaluate the truth value of the following propositions using the set (1 ,3, 5, 7)








as the domain.








1. Vx: x (x2 – 4) = 0








2. Ǝx : x + 1 > 0








3. Ǝx : x (x2








– 4) = 0

1
Expert's answer
2022-03-31T08:49:49-0400

1. x=0 or (x2-4)=0

"x \not= 0"

So it is false

2.x+1>0

1+1=2>0

So, it is true

3. x(x2-4)=0

"x=0"

"x^2-4=0"

"x= \plusmn 2"

In our set it is no 0,2,-2.

So it is false